# Moving towards informative and actionable social media research

### v1

@@ Introduction
The experiments that change some aspects of the news feed algorithm also clearly violate SUTVA as the treatment outcome depends on the status of others. To reduce co-partisan content, one study algorithmically down-ranked such content in treated users’ feeds  (?). What a user ultimately sees is, however, also dependent on a candidate set of posts shared by their connections, which is implicitly shaped by how algorithmic curation impacts their feeds. As the study was only carried out on a small portion of the U.S. population, whatever co-partisan and cross-cutting items participants saw were determined in large part by individuals who were exposed to the unmodified algorithm. This directly parallels the example of altering the load on specific power stations, without accounting for compensatory dynamics between individual units.
@@
The most recent Facebook deactivation study by Allcott et al.  (?), by contrast, exhibits inferential challenges from hysteresis. To the extent that users’ voting decisions were influenced by Facebook, opinions on the incumbent candidate (and perhaps the challenger, if well known) would have been formed over years prior to the election and experiment. There is no empirical reason to expect that withdrawal from Facebook has a symmetric effect to being exposed to the algorithm (likely for much longer time periods). The observed point-estimated magnitude of the vote shift being on the scale of what might tip an election (2.6%) is therefore striking, as even a zero percent change could be entirely consistent with preferences accrued over many years. The fact that the change was nonetheless non-significant (after adjusting for multiple comparisons) highlights an additional problem of acquiring sufficient statistical power in individual-level experiments to identify effects that are meaningful at the system level.
@@
Unactionable Research
@@
If we were to learn that social media use generally increases polarization, it could lead to the insight that decreasing or eliminating social media use might reduce affective polarization. However, social media has become digital public infrastructure with an array of effects, some good and some bad and its elimination is only thinkable for specific, vulnerable groups, such as children (a policy decision that indeed was recently made in Australia). Embracing rather than erasing heterogeneity across users is an important lever for developing interventions that can be applied to address specific harms. For instance, rather than asking whether or not social media as a whole causes decreasing well-being on average, one could instead ask if engagement-based ranking leads to unhealthy overuse.
It follows that rather than focus on the net effect, actionable research on social media should seek to identify realistic interventions concerning specific affordances of platforms that could actually be improved  (?). This could include platform interventions, such as bridging or down-ranking of toxic content, or user-side interventions such as cross-partisan dialogue  (?). Some of the Facebook studies  (?, ?) actually tested realistic interventions in the form of altered ranking algorithms, but did not assess their collective outcomes at the group-level. We encourage the research community to shift focus in this direction, in combination with the appropriate methods. In short, evaluating interventions has three key benefits over evaluating the effects of social media as a whole: it involves a coherent counterfactual and is thus a conceptually meaningful estimand; it is practically possible to study, because smaller treatments can be put in place for longer than e.g., deactivation studies; and it can guide the design of policy interventions.
@@
Adding the Complex Systems Perspective
@@
The challenges of studying social media just described are serious, and our discussion above highlights that progress forward cannot occur via solely relying on RCTs to ask unactionable questions. Fortunately, this is not uncharted territory. Conservation biologists, climate scientist, and public health researchers are rarely able to evaluate their proposed interventions using RCTs at the scale of the whole systems on which interventions are targeted. Nevertheless, they have made substantial progress in developing workable interventions for protecting ecosystems, mitigating climate change, and reducing the large-scale burden of disease. Progress in each of these areas has been heavily reliant on developing mechanistic theoretical models of the underlying processes governing the system and evaluating how interventions on that system might manifest in downstream changes. (?)
@@
As with work in climate science and conservation, the search for interventions is often motivated by observational research. Much of the research on social media is similarly motivated by observations that suggest social media is having deleterious societal impacts. Relevant outcome variables include mental health indicators, beliefs that might undermine democracy, attitudes such as affective polarization, and behaviors such as vaccination, voting, or crime. Often these variables exhibit concerning and consistent trends, such as reductions in teen mental health, rising polarization, erosion of democratic norms, and reduced rates of vaccination. So what do we do with these patterns?
@@
Too often, such observational findings are dismissed individually for heuristically lacking ”gold standard” status associated with RCTs. It is important to note, however, that that misunderstandings regarding the intrinsic causality of RCTs has a flip-side for observational work, where it is assumed they can solely produce associational or correlational evidence. Both misunderstandings arise from believing that a method imbues causality, when causality is instead a product of the assumptions we make about the associations we observe in our data both within and across studies  (?). Techniques of causal inference from observational data leverage naturally-occurring, rather than experimentally induced variation (?).
Of course the assumptions of any given observational study, which as with RCTs, can undermine inference. For example, many instrumental-variable studies in the field rely on the “exogenous” or “as-if random” roll out of infrastructure, this variation evaporates as technology proliferates and requires either explicitly accounting for non-random causes of roll out (e.g. socio-economic status, variation in jurisdiction) or assuming roll out is approximately random. Furthermore, instrumental-variable methods rely on the assumption that the instrument affects the outcome only through the variable of interest (the “treatment” variable), which may be difficult to justify in the complex and interconnected context of social media research (see here for a recent example of exclusion restriction violations  (?)). Finally, the interpretation of causal effects estimated through instrumental-variable analyses is limited to a specific subpopulation (the “compliers”), as the analyses focus on the estimation of local average treatment effects (LATE), which can considerably reduce the generalizability of the findings to the population of interest.
At first glance, it may seem that observational research, like RCTs, is limited by its assumptions that can undermine inference. Yet a key distinction is that a given line of observational evidence often has its own unique assumptions, such that consistency across approaches can provide some comfort that the phenomenon of interest is not merely an artifact of one particular assumption failure. In this sense, concerning observational patterns can be a guiding light towards domains that require some form of interventions. However, these large-scale and perhaps long-duration trends rarely provide firm, actionable mechanistic insight into design changes that inherently impact individuals on shorter time-scales.
@@
Ultimately a bridge is needed between RCTs that provides insight at the scale of individuals in discrete moments, and observational approaches that reveal trends at larger time-scales. Other complexity minded disciplines have built such bridges using an empirically-grounded theory- and model-driven paradigm that aims at understanding specific mechanisms. Such an approach of triangulating experimental, observational and theoretical work (Fig. 2), which we feel is necessary for the field to move forward, stands in contrast to relying on single-shot experiments to generate broad statements about platforms’ overall impacts. Although uncommon in the social sciences, this paradigm is not out of reach. Theory and mechanistic models do not serve to replace empirical research, but rather augment it by bridging gaps across scales of organizational complexity – from individual behavior to societal-level phenomena. If observational approaches to causal inference reveal a trend, theoretical approaches can be used to propose how such trends might arise from mechanisms at play. In turn, the mechanistic assumptions of the multi-scale model can be identified and confirmed at individual scales with formal experiments, providing confidence that the model is able to capture leading-order features governing dynamics across scales. Computational methods, such as agent-based simulations or interactive online experiments furthermore allow for bridging the methodological gaps.
@@
The spread of misleading information online provides a salient example of how triangulation across observational and experimental research can be facilitated by complexity-minded theory. Given observations that misinformation proliferates online, a variety of interventions have been proposed, studied, experimentally validated in RCTs  (?). However, despite the success in RCTs and even convincing some platforms to adopt these approaches, estimating their population-level effect remains is not straightforward as the dynamics of misinformation spreading present a typical example of a complex, non-linear system  (?).
@@
Yet misinformation spreads in a manner that can be captured by models of contagious processes  (?). Beyond mere theory, data can be used to constrain parameters of contagion models, providing a tight fit to observed dynamics and allowing for evaluation of counterfactuals in which various policies are implemented  (?, ?). Contagion models can be explicitly made consistent with evidence from RCTs, by hard-coding the observed reductions in sharing behavior from experiments. Using such a model fit to data from the 2020 US elections, it was predicted that despite the nominally large effect size of the “community notes” program which relies on appending crowd-sourced fact-checks  (?), changes in exposure would be minimal—as large amounts of spread by influential accounts can occur prior to labeling and for unlabeled posts. Observational investigations provided confirmation of these model implications, confirming that even as community notes were indeed too slow to significantly reduce engagement with misinformation  (?).
@@
These mechanistic models can often be used to evaluate multiple potential interventions in a common framework. Here, for example, the model indicates that dynamics of information spread on platforms will be highly determined by the actions of influential accounts  (?). While content creators have disproportionate agency in shaping the content that is created and propagated, which is sometimes cited as an argument against the relative importance of social media platforms, a complex systems view recognizes their role in dependence of specific incentive structures that are set by the platforms  (?). For example, by using economic models, as content creation on platforms is labor  (?). And this suggestion that influence incentives provide a uniquely large lever for intervention is indicated on longer time-horizons by observational research by Frimer et al., which measured a 23% increase in incivility among US politicians and attributed it partly to learning from positive feedback for uncivil tweets  (?). As interventions at the creator level have been comparatively understudied relative to user-specific interventions, this suggests a fruitful avenue for future research. For example, we might ask what can break feedback loops between creators and their audiences that may be driving influencers towards extremism.
@@
The critical thing is that we can either, as a field, opt to continue scratching our heads over conflicting evidence or attempt to bring it all together into a common framework. Recognizing large-scale patterns that are consistent across observational approaches which vary in their assumptions can guide us towards where interventions may be needed. RCTs can help us understand proximate, small-scale effects that in turn can fuel theory linking observations and intervention. In triangulating across these modalities of inference, research on social media can follow in the footsteps of other disciplines that have embraced complexity to develop workable interventions in complex systems (Fig. 2.
@@ Discussion
A considerable body of research on the causal impacts of social media has chased an unanswerable, and unactionable question of whether the world would be, or would have been, better off without social media. Disproportionate weight has been given to RCTs which, in failing to account for complexity, cannot yield the desired insight into impacts of platforms at societal scales. Trying to reconcile evidence from RCTs with either proximal patterns of content exposure or more distal patterns that emerge at societal scales, such as polarization, is a recipe for failing to reach consensus and inaction.
@@
More generally, it is worth reflecting on our inferential goals when establishing links between platforms and harms. Such quantities are likely to be time-varying, hysteretic, and heterogeneous such that it is unlikely that we could come up with a single estimate of effects on any given domain of concern—much less across domains. Even if we could, such an estimate could do little more than motivate an attempt to extricate social media from society—an intervention that is difficult to execute and could very well do more harm than good. Instead, the key value in any such apparent causal links, however tenuous, is a recognition that social media can impact large-scale phenomena of interest. This, in turn, motivates the search for design principles and regulatory action that can reduce risks and enhance benefits.
@@
Perhaps the first question we can ask is whether we need a causal understanding of larger-scale effects to warrant intervention. Normative and legal justifications for altering the distribution of content may in many cases be sufficient. Researchers may be able to best support such interventions by simply documenting failures of platforms preventing the distribution of illegal content of their own accord. The Digital Services Act of the European Commission is a step towards enabling researchers to do so. Platform interventions could go beyond content moderation and build on our existing understanding of how algorithmic design choice shapes content amplification. At present these choices are decided based on business priorities, and there is little reason they should not, or cannot reflect widely held prosocial goals. Moreover, as the content shared on a platform is disproportionately shaped by the incentives that govern content creation, understanding the behavior and dynamics of content creators is likely to be a particularly productive line of research.
@@
When more firm causal links across scales of organizational complexity are required, it is important to recognize that no approach in social media research represents the “gold standard” possible in other scientific domains. As a result, we should not anticipate that any given modality of inference (e.g., RCTs, observation, theory) will be sufficient to identify phenomena and propose interventions. Instead, progress will require embracing complexity and relying on a triangulation across lines of evidence with varying assumption, leveraging computational and mathematical models to link and infer dynamics across time and scale.
@@
We believe this is not only a possible path forward for the field, but a necessary one. The alternative is to endlessly fail to reconcile conflicting evidence, chasing the ghosts of overlooked assumptions and unanswerable questions. Doing so not only fails to yield actionable insight, but undermines any motivations by stakeholders to search for interventions in the first place. Given this, the community should set aside notions of rote causal impacts and move towards more actionable, complexity minded, and informative social media research.
@@
  • Competing interests. SL also receives funding from Jigsaw (a technology incubator created by Google). SL, PLS and AO also work with the Joint Research Centre of the European Union, both in a paid and unpaid capacity. J.B.-C. has engaged in paid consulting for the United Nations.

### v2

@@ Introduction
The experiments that change some aspects of the news feed algorithm also clearly violate SUTVA as the treatment outcome depends on the status of others. To reduce co-partisan content, one study algorithmically down-ranked such content in treated users’ feeds (?). What a user ultimately sees is, however, also dependent on a candidate set of posts shared by their connections, which is implicitly shaped by how algorithmic curation impacts their feeds. As the study was only carried out on a small portion of the U.S. population, whatever co-partisan and cross-cutting items participants saw were determined in large part by individuals who were exposed to the unmodified algorithm. As a result, even the individual-level effects in this study are likely biased towards zero by unaddressed SUTVA violations.
@@
The most recent Facebook deactivation study by Allcott et al. (?), by contrast, exhibits inferential challenges from hysteresis. A key outcome of interest was whether Facebook had any impact on election outcomes in the U.S. Yet to the extent that users’ voting decisions were influenced by Facebook, opinions on the incumbent candidate (and perhaps the challenger, if well known) would have been formed over years prior to the election and experiment. There is no empirical reason to expect that withdrawal from Facebook—as with Duolingo in our earlier example—has a symmetric effect to being exposed to the algorithm (likely for much longer time periods). The observed point-estimated magnitude of the vote shift being on the scale of what might tip an election (2.6%) is therefore striking, as even a zero percent change could be entirely consistent with preferences accrued over many years. As the observed vote share change is almost certainly an underestimate of the total causal effect, as in the case of early research on voter turnout described above, the total impact of Meta on elections may well have been appreciable or even determinative.
@@
Our critique of RCTs and the Meta collaboration is not to say that such approaches are without value in the study of social media, and we return to their potential later. However, it is important to move past the notion that RCTs inherently and intrinsically produce unbiased causal estimates. Instead, the quality of any causal insight–=from experiments to observational data analysis—depends wholly on the assumptions underlying the causal inference and not the intrinsic properties of the approach.
As we have outlined above, the assumptions required by RCTs are by no means guaranteed in the study of social media. Unless the outcome of interest is on the individual level and the treatment independent of others (e.g., attention allocation as a function of linguistic features of a post), estimates of individual-level effects may not provide reliable insight into societal phenomena, and may themselves be under-estimated if issues such as SUTVA go unaddressed. Taken together, claims that small, varying, or null effects from RCTs on the individual level alleviate concerns about broader societal impacts of social media have to be critically examined.
From causal effects to actionable understanding
@@
If we were to learn that social media use generally increases polarization, it could lead to the insight that decreasing or eliminating social media use might reduce affective polarization. However, social media has become digital public infrastructure with an array of effects, some good and some bad. Its elimination is only within range of political possibility for specific, vulnerable groups, such as children (a policy decision that indeed was recently made in Australia). Examining, rather than erasing heterogeneity across users, as well as testing feasible interventions are important levers for developing policies that can be applied to address specific harms (?). For instance, rather than asking whether or not social media as a whole causes decreasing well-being on average, one could instead ask if engagement-based ranking leads to unhealthy overuse among teenagers, if one is concerned about real-word implementation of the findings.
It follows that rather than focus on the net effect, actionable research on social media should seek to identify realistic interventions concerning specific affordances of platforms that could actually be improved (?). This could include platform interventions, such as bridging or down-ranking of toxic content, or user-side interventions such as cross-partisan dialogue (?). Some of the Meta studies discussed earlier (?, ?) actually tested realistic interventions in the form of altered ranking algorithms, but did not assess their collective outcomes at the group-level, which almost certainly lead to under-estimation of their causal potential. Had they been done differently, in particular without violating SUTVA, they may have yielded more valuable outcomes.
We encourage the research community to shift focus in this direction, in combination with the appropriate methods. In short, evaluating interventions on a group-level has three key benefits over evaluating the effects of social media as a whole: it involves a coherent counterfactual and is thus a conceptually meaningful estimand; it is practically possible to study, because smaller treatments can be put in place for longer than e.g., deactivation studies; and it can guide the design of policy interventions.
@@
Embracing Epistemic Pluralism
@@
The challenges of studying social media just described are serious, and our discussion above highlights that progress cannot occur by solely relying on RCTs, and their findings should not implicitly take causal precedence over other modes of inference. Fortunately, this is not uncharted territory. Conservation biologists, climate scientists, and public health researchers are rarely able to evaluate their proposed interventions using RCTs at the scale of the whole systems on which interventions are targeted. Nevertheless, they have made substantial progress in developing workable interventions for protecting ecosystems, mitigating climate change, and reducing the large-scale burden of disease. Progress in each of these areas has been heavily reliant on epistemically pluralistic approaches the leverage the strengths and weaknesses of experiments, theory, and observation.
@@
As with work in climate science and conservation, the search for interventions is often motivated by observational research. Much of the research on social media is similarly motivated by observations that suggest social media is having deleterious societal impacts. Relevant outcome variables include mental health indicators, beliefs that might undermine democracy, attitudes such as affective polarization, and behaviors such as vaccination, voting, or hate crime. Often these variables exhibit concerning and consistent trends, such as reductions in teen mental health, rising polarization, erosion of democratic norms, and reduced rates of vaccination. So what do we do with these patterns?
@@
Such observational findings are often dismissed for lacking the “gold standard” status thought to be associated with RCTs. This is, however, merely an inverse of the misunderstanding about the intrinsic causality of RCTs. The assumption that observational studies can solely produce associational or correlational evidence is as fraught as the assumption that RCTs necessarily yield causal insights. Both misunderstandings arise from believing that a method imbues causality, whereas causality is actually a product of the assumptions we make about the associations we observe in our data both within and across studies (?). For that reason, observational data can certainly identify causal relationships (?).
Of course the assumptions of any given observational study, as with RCTs, can either undermine or buttress causal inference. We introduced the idea of instrumental variables and related techniques earlier. Causal inference in these contexts makes assumptions that instruments used, such as infrastructure roll-outs or natural disasters that divert media attention, are “exogenous” or “as-if random”.
Instrumental-variable methods additionally rely on the assumption that the instrument affects the outcome only through the variable of interest (the “treatment” variable, e.g., media coverage), which may occasionally be difficult to justify in the complex and interconnected context of social media research (see here for a recent example of exclusion restriction violations (?)). This often requires adding co-variates to the statistical model to adjust for sources of bias and ideally identify the causal pathway of interest (?, ?, ?). However simply adding a large suite of co-variates to the model can just as easily introduce or amplify bias as remove it, such as co-variates measured post-treatment or colliders (?). Yet, even with residual confounding or bias techniques exist for quantifying or constraining their magnitude (?). Altogether, whether observational data represents mere associations or causality ultimately rests on the degree to which an explicit causal model’s assumptions are reasonably justified.
At first glance, it may seem that observational research, similar to RCTs, is limited by assumptions that can undermine any attempted inference. There is however a crucial difference: different observational approaches come with their own set of assumptions. If results turn out to be consistent across approaches despite the variability in assumptions, this increases confidence that the effect of interest is not merely an artifact of one particular assumption failure. Observational patterns that reveal concerning impacts of social media, even if inconclusive in isolation, can therefore be a guiding light towards specific domains that require some form of interventions.
@@
This perspective on causality allows us to recover the role of RCTs in social media research. For some contexts, questions, and inferential goals their assumptions will hold reasonably well and they can be powerful tools for insight. Consider common experiments that evaluate changes in mental health or academic performance during a multi-week period of giving up social media. While this cannot tell us about the societal effect of social media, the estimate it produces could certainly inform an individual’s decision to take a break. SUTVA violations here are less of a concern as the person taking a break—as with the participants—remain in a world filled with social media. In other contexts, we may be able to make progress using RCTs and reasonable assumptions about how their effects scale. For example, a design intervention that reduces violent content in one’s news feed might support a reasonable inference that applying it broadly would reduce the reach of violent content on platform.
More generally, RCTs can be remarkably useful provided we are clear on the assumptions they require, that we are making, and how each relate to our inferential goals. While technically difficult and expensive, RCTs that change the unit of randomization from individual to a group-size of interest (e.g., online communities) become a powerful tool for answering causal, “what if” questions for social media design (?, ?).
@@
Ultimately a bridge is needed between RCTs that provide insight at the scale of individuals in discrete moments, and observational approaches that reveal trends at larger time-scales. Other disciplines tasked with understanding and intervening on complex systems have built such bridges using an empirically-grounded theory- and model-driven paradigm that aims at understanding specific mechanisms. Such an approach of bridging experimental, observational and theoretical work (Fig. 2), which we feel is necessary for the field to move forward, stands in contrast to relying on single-shot experiments to generate broad statements about platforms’ overall impacts.
Although uncommon in the social sciences, this paradigm is not out of reach. Theory and mechanistic models do not serve to replace empirical research, but rather augment it by bridging gaps across scales of organizational complexity—from individual behavior to societal-level phenomena. If observational approaches to causal inference reveal a trend, theoretical approaches can be used to propose how such trends might arise from mechanisms at play. In turn, the mechanistic assumptions of the multi-scale model can be identified and confirmed at individual scales with formal experiments, providing confidence that the model is able to capture leading-order features governing dynamics across scales. Computational methods, such as agent-based simulations or interactive online experiments furthermore allow for bridging the methodological gaps (?).
@@
The spread of misleading information online provides a salient example of how jointly leveraging observational and experimental research can be facilitated by complexity-minded theory. Given observations that misinformation proliferates online, a variety of interventions have been proposed, studied, and experimentally validated in RCTs (?). However, despite the success in RCTs and even convincing some platforms to adopt these approaches, estimating their population-level effect is not straightforward as the dynamics of misinformation spreading present a typical example of a complex, non-linear system (?).
@@
Yet misinformation spreads in a manner that can be captured by models of contagious processes (?). Beyond mere theory, data can be used to constrain parameters of contagion models, providing a tight fit to observed dynamics and allowing for evaluation of counterfactuals in which various policies are implemented (?, ?). Contagion models can be explicitly made consistent with evidence from RCTs, by hard-coding the observed reductions in sharing behavior from experiments.
Using such a model fit to data from the 2020 US elections, it was predicted that despite the nominally large effect size of the “community notes” program which relies on appending crowd-sourced fact-checks (?), changes in exposure would be minimal—as large amounts of information dissemination by influential accounts can occur prior to labeling and for unlabeled posts. Observational investigations provided confirmation of these model implications, confirming that community notes were indeed too slow to significantly reduce engagement with misinformation (?).
@@
These mechanistic models can often be used to evaluate multiple potential interventions in a common framework. Here, for example, the model indicates that dynamics of information spread on platforms will be highly determined by the actions of influential accounts (?). While content creators have disproportionate agency in shaping the content that is created and propagated, which is sometimes cited as an argument against the relative importance of social media platforms, a broader perspective recognizes their role in dependence of specific incentive structures that are set by the platforms (?). For example, by using economic models, as content creation on platforms is labor (?). And this suggestion that influence incentives provide a uniquely large lever for intervention is indicated on longer time-horizons by observational research by Frimer et al., which measured a 23% increase in incivility among US politicians and attributed it partly to learning from positive feedback for uncivil tweets (?). As interventions at the creator level have been comparatively understudied relative to user-specific interventions, this suggests a fruitful avenue for future research. For example, we might ask what can break feedback loops between creators and their audiences that may be driving influencers towards extremism.
@@
The critical issue is that we can either, as a field, opt to continue scratching our heads over conflicting evidence or attempt to bring it all together into a common framework. Recognizing large-scale patterns that are consistent across observational approaches which vary in their assumptions can guide us towards where interventions may be needed. RCTs can help us understand proximate, small-scale effects that in turn can fuel theory linking observations and intervention. In embracing multiple modalities of inference, research on social media can follow in the footsteps of other disciplines that have embraced complexity to develop workable interventions in complex systems (Fig. 2).
@@ Discussion
Summaries of social media’s impacts have tended to overlook the affordances of observational research while failing to examine the limitations of RCTs—asserting erroneously that evidence from the latter outweighs the former (?, ?). In doing so, it has been easy to identify concerns yet difficult to reach consensus on societal effects. Actionable interventions at the scale of the problem have remained elusive.
@@
More generally, it is worth reflecting on our inferential goals when establishing links between platforms and harms. Such quantities are likely to be time-varying, hysteretic, and heterogeneous such that it is unlikely that we could come up with a single estimate of effects on any given domain of concern—much less across domains. Beyond debates about harm, benefits, or lack thereof, the key value in identifying apparent causal links, however tenuous, is a recognition that social media has the potential to impact large-scale phenomena of interest. This, in turn, motivates the search for design principles and regulatory action that can reduce risks and enhance benefits.
@@
Perhaps the first question we can ask is whether causality is required to warrant intervention. Normative and legal justifications for altering the distribution of content may in many cases be sufficient. Researchers may be able to best support such interventions by simply documenting failures of platforms preventing the distribution of illegal content of their own accord. The Digital Services Act of the European Union is a step towards enabling researchers to do so by providing an access mechanism for platform data (Article 40). Platform interventions could go beyond content moderation and build on our existing understanding of how algorithmic design choice shapes content amplification. At present these choices are decided based on business priorities, and there is little reason they should not, or cannot reflect widely held prosocial goals. Moreover, as the content shared on a platform is disproportionately shaped by the incentives that govern content creation, understanding the behavior and dynamics of content creators is likely to be a particularly productive line of research.
@@
When more firm causal links across scales of organizational complexity are required, it is important to recognize that no approach in social media research represents the “gold standard” possible in other scientific domains. As a result, we should not anticipate that any given modality of inference (e.g., RCTs, observation, theory) will be sufficient to identify phenomena and propose interventions. Instead, progress will require embracing complexity and relying on epistemic plurality across lines of evidence with varying assumptions, leveraging computational and mathematical models to link and infer dynamics across time and scale.
@@
We believe this is not only a possible path forward for the field, but a necessary one. The alternative is to endlessly fail to reconcile conflicting evidence, chasing the ghosts of overlooked assumptions and unanswerable questions. Doing so not only fails to yield actionable insight, but undermines any motivations by stakeholders to search for interventions in the first place. Given this, the community should set aside notions of rote causal impacts and move towards more actionable, epistemically pluralistic, and informative social media research.
@@
  • Competing interests. SL also receives funding from Jigsaw (a technology incubator created by Google). SL, PLS and AO also work with the Joint Research Centre of the European Union, both in a paid and unpaid capacity. J.B.-C. has engaged in paid consulting for the United Nations. A.N. has engaged in paid consulting in litigation regarding social media and adolescent mental health.